Gear Drive motor Input A B D E Shaft Gear R - Gear S Output to crusher sis is etis olis Output to conveyor Gear C Gear A 101.q=154
Answer #1

Solution:- = 127 + 154 127-154 784a7 1547 = 127: 154 1187,154 127.154 Geanta- 100 Powevy p= 20+ - 80+ 100 diameter of gear daTangential Force FLE = TE 919721 6920:34N viadial fource fire = fte Tand 2518.79 N = 6920.34 x Tm (20) = Geauto:- shaft is noResolving vertically- A 197-164 B 1937.1545 127154 D 124-154 € FLA RON ELE Rod Root Roy+FEA +fte = ftc Rort RO - 4132.8 - 807Resolving horizontalya ROH fore Fue - Fora Rosh RoH ROHT ROH + 2938.41 + 25 18.79 = 1504.2 ROH + ROH = 1504.2-54572 = -3953 NResulbing Bending Moment: Me = faulot (Maher (1026542)*7(8+3630-55 = 1090871-21 = [1090871 Nm] Me= coulant (maler (1214357) +